OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. S(1/2, 1), T(1/2, 4)

OpenStudy (aum):

First find the slope from those two points. Slope \(\Large m = \frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

not sure how to do that with the 1/2

OpenStudy (larseighner):

Treat it like 0.5 (that is what it is)

OpenStudy (larseighner):

If you are not using a calculator, you can multiply the whole fraction by 2/2. That gets rid of the halves, and multiplying by 1 won't change anything.

OpenStudy (aum):

Well, in this case, the x value happens to be the same for both points. Therefore, there is no need to even calculate slope. If x value is the same it means the line is a vertical line, parallel to the y-axis. (If y value is the same it means the line is a horizontal line, parallel to the x-axis.)

OpenStudy (anonymous):

-2x + y = 0 is this right

OpenStudy (aum):

Also, if slope is infinity it means the line is a vertical line.

OpenStudy (larseighner):

No. As @aum points out, you cannot figure slope using the two points because \(x_1 - x_0 = 0\) and division by zero is a no, no.

OpenStudy (aum):

The equation for a vertical line is x = constant (you will have to find the constant). In this case x is always 1/2. Therefore, the equation of the line is x = 1/2

OpenStudy (anonymous):

thank you I hope I get this soon its hard

OpenStudy (aum):

It will get easier with practice.

OpenStudy (anonymous):

can you please help with one more

OpenStudy (aum):

go ahead

OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (aum):

The point-slope form of a line is: \(\Large y-y_1 = m(x-x_1)\) where m is the slope and \(x_1 = 5, y_1 = -2\)

OpenStudy (anonymous):

so how do I do that then I am so lost

OpenStudy (aum):

\(\Large y-y_1 = m(x-x_1)\) Put \(\Large m = \frac 25, x_1 = 5, y_1 = -2\) in the above equation. Simplify. Then rearrange it to put it in the standard form.

OpenStudy (anonymous):

this is hard not having no fun thank you I will keep tring to get it with this I am sure I will sooner or later thank you so much for helping :)

OpenStudy (aum):

y - y1 = m(x - x1) m = 2/5, x1 = 5, y1 = -2 y - (-2) = 2/5 * (x - 5) y + 2 = 2/5 * (x - 5) To get rid of the fraction on the right multiply both sides by 5: 5y + 10 = 2 * (x-5) 5y + 10 = 2x - 10 subtract 10 from both sides 5y = 2x - 20 Put this in standard form which is Ax + By = C subtract 5y from both sides: 0 = 2x - 5y - 20 add 20 to both sides: 20 = 2x - 5y or 2x - 5y = 20